# On the Performance of the Neyman Allocation with Small Pilots

### v3

@@ Appendix B Additional Empirical Examples
B.4 [16, ]
Partnering with the Peruvian nongovernmental organization PRISMA, [16, ] conducted two RCTs to investigate the efficacy of various interventions in encouraging recycling. First, the Participation Study considers the following $9$ different messaging strategies and their relative success in enrolling members into recycling programs:
@@
    “Participates any time” is an indicator that takes the value 1 if a household turned in residuals over the course of the study.
@@
    “Participation Ratio” is the number of times a household turns in residual over the total number of opportunities they had to turn in residuals.
@@
    “Participates during either of last two visits” is an indicator that takes value 1 if the household turned in residual during one of the last two canvassing weeks.
The results in Table 3 of [16, ] shows that messaging had no effect in increasing participation in the program. Table 13 displays the standard deviation of the various outcomes by treatment type. Table 14, shows the ratio of the standard deviation in the outcome variable of each treatment group, with respect to that of the control group. Here, we see that the outcomes are highly homoskedastic, suggesting little scope for improvement over the naive experiment.
@@
B.5 [20, ]
[20, ] study employers’ perceptions of the value of post-secondary degrees using a field experiment. The experimental units are fictitious resumes to be used in applications to vacancies posted on a large online job board. Their focus is on degrees and certificates awarded in the two largest occupational categories in the United States: business and health. Resumes are randomly assigned sector and selectivity of (degree-awarding) institutions. Fictitious resumes are created using a vast online database of actual resumes of job seekers, with applicant characteristics varying randomly (i.e. characteristics are randomly assigned). Outcomes are callback rates. There are three main comparisons in the paper:
@@
[20, ] find that BA degrees in business from large online for-profit institutions are more 22% less likely to receive a callback than applicants with similar degrees from non-selective public schools when the job vacancy requires a BA. When a business job opening does not list a BA requirement, they find no significant overall advantage to having a post-secondary degree. For health jobs, resumes with certificates from for-profit institutions are 57% less likely to receive a callback than those with similar certificates from public institutions when the job listing does not require a post-secondary certificate. No significant difference in callback rates are found when the health job listing requires a certificate.
Table 16 reports standard deviations across treatment arms across the various sub-populations of interest in [20, ]. Since in most sub-populations, there are more than two treatment arms, we do not report ratios. However, pairwise comparisons between treatment arms within any chosen subpopulation shows strong evidence of relative homoskedasticity across the board.
@@
B.6 [7, ]
[7, ] study the effect of working from home on employees’ productivity via a randomized experiment at Ctrip, a NASDAQ-listed Chinese travel agency with 16000 employees. The main concern is whether or not working from leads to shirking. Ctrip decided to run a nine-month experiment on working from home. They asked the 996 employees in the airfare and hotel departments of the Shanghai call center if they were interested in working from home four days a week and one day in the office. 503 of these employees were interested and of these, 249 were qualified to take part on the basis of tenure, broadband access and access to private work space at home. Qualified employees were assigned to working from home if they had even-numbered birthdays so that those with odd-numbered birthdays formed the control group. The treatment and control groups were comprised of 131 and 118 employees respectively. The only difference between the two groups was location of work – both groups used the same equipment, faced the same workload and were compensated under the same pay system. The authors find a 13% increase in productivity of which the main source of improvement was a 9% increase in the number of minutes worked during a shift. The remaining 4% came from an increase in the number of calls per minute worked.
Table 17 reports standard deviations for treatment and control groups in the experiment as well as the standard deviation ratios for the main outcomes of interest in [7, ]. Strong evidence of relative homoskedasticity with respect to treatment status presents in this study as well – suggesting that the balanced allocation would outperform the FNA.
@@
B.7 [23, ]
[23, ] use the Honduran PRAF experiment to study the impact of conditional cash transfers (CCT) on the likelihood of children to work versus enrolling in school. The PRAF experiment randomly allocated CCT’s among 70 municipalities. These 70 were chosen out of a total of 298 on the basis of mean heights-for-age z-scores of first graders. The 70 municipalities were further assigned to four treatment arms termed G1, G2, G3, G4. G1 received CCT’s in education and health. G2 received CCT’s in addition to direct investment in education and health centers. G3 received only direct investments and finally, G4 served as the control group and received no interventions. The 70 municipalities were further divided into 5 strata each consisting of 14 municipalities on the basis of quintiles of mean height-for-age. Random assignment was performed within these strata (stratified randomization). The final sample consisted of 20 municipalities in G1, 20 in G2, 10 in G3, and 20 in G4. The authors match the experimental data with census data and use the latter to construct the outcomes of interest which are three dummy variables. The first is an indicator for whether a child is enrolled in and attending school during the time of the census. The second indicates whether the child worked during the week prior to the census or conditional on a negative response to the former, whether they reported non-wage employment in a family farm or business. The third indicates whether the child worked exclusively on household chores. The authors find that overall, children eligible for CCTs were 8% more likely to enroll in school and 3% less likely to work.
@@
B.8 [10, ]
[10, ] conduct a field experiment to study efficacy of peer intermediation in mitigating adverse selection and moral hazard in credit markets. To identify the effects of peer screening and enforcement, they use a two-stage referral incentive field experiment. The experiment was conducted through Opportunity Finance South Africa (Opportunity), a for-profit lender in the consumer micro-loan market. Over the period of February 2008 through July 2009, Opportunity offered individuals approved for a loan the option to participate in its “Refer-A-Friend” program. Referred individuals earned R40 if they brought in a referral card and were approved for a loan. The referrer could earn R100 for referring someone who was subsequently approved for and/or repaid the loan, depending on the referrer’s incentive contract. Referrers were randomly assigned to one of two ex-ante incentive contracts:
@@
Table 19 reports standard deviations in the main outcomes of interest in [10, ] across the four treatment arms. For a given outcome, pairwise comparisons across arms yields evidence of relative homoskedasticity with respect to treatment status. We conclude that in this case, the balanced allocation would outperform the FNA.

### v4

@@ B.3 McKenzie (2017)
B.4 Chong et al. (2015)
Partnering with the Peruvian nongovernmental organization PRISMA, Chong et al. (2015) conducted two RCTs to investigate the efficacy of various interventions in encouraging recycling. First, the Participation Study considers the following $9$ different messaging strategies and their relative success in enrolling members into recycling programs:
@@
    “Participates any time" is an indicator that takes the value 1 if a household turned in residuals over the course of the study.
@@
    “Participation Ratio" is the number of times a household turns in residual over the total number of opportunities they had to turn in residuals.
@@
    “Participates during either of last two visits" is an indicator that takes value 1 if the household turned in residual during one of the last two canvassing weeks.
The results in Table 3 of Chong et al. (2015) shows that messaging had no effect in increasing participation in the program. Table 13 displays the standard deviation of the various outcomes by treatment type. Table 14, shows the ratio of the standard deviation in the outcome variable of each treatment group, with respect to that of the control group. Here, we see that the outcomes are highly homoskedastic, suggesting little scope for improvement over the naive experiment.
@@
B.5 Deming et al. (2016)
Deming et al. (2016) study employers’ perceptions of the value of post-secondary degrees using a field experiment. The experimental units are fictitious resumes to be used in applications to vacancies posted on a large online job board. Their focus is on degrees and certificates awarded in the two largest occupational categories in the United States: business and health. Resumes are randomly assigned sector and selectivity of (degree-awarding) institutions. Fictitious resumes are created using a vast online database of actual resumes of job seekers, with applicant characteristics varying randomly (i.e. characteristics are randomly assigned). Outcomes are callback rates. There are three main comparisons in the paper:
@@
Deming et al. (2016) find that BA degrees in business from large online for-profit institutions are more 22% less likely to receive a callback than applicants with similar degrees from non-selective public schools when the job vacancy requires a BA. When a business job opening does not list a BA requirement, they find no significant overall advantage to having a post-secondary degree. For health jobs, resumes with certificates from for-profit institutions are 57% less likely to receive a callback than those with similar certificates from public institutions when the job listing does not require a post-secondary certificate. No significant difference in callback rates are found when the health job listing requires a certificate.
Table 16 reports standard deviations across treatment arms across the various sub-populations of interest in Deming et al. (2016). Since in most sub-populations, there are more than two treatment arms, we do not report ratios. However, pairwise comparisons between treatment arms within any chosen subpopulation shows strong evidence of relative homoskedasticity across the board.
@@
B.6 Bloom et al. (2014)
Bloom et al. (2014) study the effect of working from home on employees’ productivity via a randomized experiment at Ctrip, a NASDAQ-listed Chinese travel agency with 16000 employees. The main concern is whether or not working from leads to shirking. Ctrip decided to run a nine-month experiment on working from home. They asked the 996 employees in the airfare and hotel departments of the Shanghai call center if they were interested in working from home four days a week and one day in the office. 503 of these employees were interested and of these, 249 were qualified to take part on the basis of tenure, broadband access and access to private work space at home. Qualified employees were assigned to working from home if they had even-numbered birthdays so that those with odd-numbered birthdays formed the control group. The treatment and control groups were comprised of 131 and 118 employees respectively. The only difference between the two groups was location of work – both groups used the same equipment, faced the same workload and were compensated under the same pay system. The authors find a 13% increase in productivity of which the main source of improvement was a 9% increase in the number of minutes worked during a shift. The remaining 4% came from an increase in the number of calls per minute worked.
Table 17 reports standard deviations for treatment and control groups in the experiment as well as the standard deviation ratios for the main outcomes of interest in Bloom et al. (2014). Strong evidence of relative homoskedasticity with respect to treatment status presents in this study as well – suggesting that the balanced allocation would outperform the FNA.
@@
B.7 Galiani and McEwan (2013)
Galiani and McEwan (2013) use the Honduran PRAF experiment to study the impact of conditional cash transfers (CCT) on the likelihood of children to work versus enrolling in school. The PRAF experiment randomly allocated CCT’s among 70 municipalities. These 70 were chosen out of a total of 298 on the basis of mean heights-for-age z-scores of first graders. The 70 municipalities were further assigned to four treatment arms termed G1, G2, G3, G4. G1 received CCT’s in education and health. G2 received CCT’s in addition to direct investment in education and health centers. G3 received only direct investments and finally, G4 served as the control group and received no interventions. The 70 municipalities were further divided into 5 strata each consisting of 14 municipalities on the basis of quintiles of mean height-for-age. Random assignment was performed within these strata (stratified randomization). The final sample consisted of 20 municipalities in G1, 20 in G2, 10 in G3, and 20 in G4. The authors match the experimental data with census data and use the latter to construct the outcomes of interest which are three dummy variables. The first is an indicator for whether a child is enrolled in and attending school during the time of the census. The second indicates whether the child worked during the week prior to the census or conditional on a negative response to the former, whether they reported non-wage employment in a family farm or business. The third indicates whether the child worked exclusively on household chores. The authors find that overall, children eligible for CCTs were 8% more likely to enroll in school and 3% less likely to work.
@@ B.7 Galiani and McEwan (2013)
B.8 Bryan et al. (2015)
Bryan et al. (2015) conduct a field experiment to study efficacy of peer intermediation in mitigating adverse selection and moral hazard in credit markets. To identify the effects of peer screening and enforcement, they use a two-stage referral incentive field experiment. The experiment was conducted through Opportunity Finance South Africa (Opportunity), a for-profit lender in the consumer micro-loan market. Over the period of February 2008 through July 2009, Opportunity offered individuals approved for a loan the option to participate in its “Refer-A-Friend” program. Referred individuals earned R40 if they brought in a referral card and were approved for a loan. The referrer could earn R100 for referring someone who was subsequently approved for and/or repaid the loan, depending on the referrer’s incentive contract. Referrers were randomly assigned to one of two ex-ante incentive contracts:
@@
Table 19 reports standard deviations in the main outcomes of interest in Bryan et al. (2015) across the four treatment arms. For a given outcome, pairwise comparisons across arms yields evidence of relative homoskedasticity with respect to treatment status. We conclude that in this case, the balanced allocation would outperform the FNA.